The density of aluminum is 2.70 g*cm^3. What is the volume ( in cm^3) and edge length of a cube of Al with a mass of 10.0 g?
1
Expert's answer
2017-11-22T13:55:07-0500
m = V * rho,
V = m / rho = 10 / 2.7 = 3.70 (cm3)
V = a3,
a = 1.54 (cm)

Answer
V = 3.7 cm3
a = 1.54 cm